OpenStudy (anonymous):

(x^2+8x+13) divide (x+5) Answer is quotient and remainder. Help please! I've given my best tries to the others...but I can't get this one...

OpenStudy (whpalmer4):

take the first term of the dividend (the thing being divided), and divide it by the first term of the divisor (the thing doing the dividing), what do you get?

OpenStudy (anonymous):

5x^2? how do you divide an x? an x with an exponent

OpenStudy (whpalmer4):

\[\frac{x^2}{x} = \]

OpenStudy (anonymous):

2x^2? or just x?

OpenStudy (whpalmer4):

pick one. then test your answer by multiplying it by \(x\) and see if you get \(x^2\) as the result. \[\frac{a}{b} =c\] only if \[c*b = a\]

OpenStudy (anonymous):

x works!

OpenStudy (anonymous):

yay learning :)

OpenStudy (whpalmer4):

very good. \[x^2 = x*x\]\[\frac{x^2}{x} =\frac{x*x}{x} = \frac{\cancel{x*}x}{\cancel{x}} = x\] Or \[\frac{x^2}{x} = \frac{x^2}{x^1} = x^{2-1} = x^1 = x\]

OpenStudy (whpalmer4):

okay, so we do this like long division...we take the first term of the thing being divided, and divide it by the first term of the thing doing the dividing. we did that, and we got \(x\) as the first term of our quotient. Now, we multiply \(x\) with the entire divisor (the thing doing the dividing), and subtract the resulting quantity from what we started with: \[x(x+5) =\]

OpenStudy (anonymous):

x^2 + 5x

OpenStudy (whpalmer4):

good. what is \[x^2+8x+13 - (x^2+5x) = \]

OpenStudy (anonymous):

do I simplify the first part first? like have 8x^3 + 13 - (blah blah)

OpenStudy (anonymous):

that seems like the thing I should do?

OpenStudy (anonymous):

am I wrong?

OpenStudy (whpalmer4):

Uh, could you simplify my expression again?

OpenStudy (anonymous):

hmmm? what do you mean?

OpenStudy (anonymous):

8x^3 + 13 and then I multiply across right?

OpenStudy (anonymous):

is your internet connection bad?

OpenStudy (whpalmer4):

no, I'm just stunned at some of the things I'm seeing :-)

OpenStudy (whpalmer4):

\[x^2+8x+13 - (x^2+5x) = x^2+8x+13-1(x^2+5x)\]Could you use the distributive property on the stuff in the parentheses, please?

OpenStudy (whpalmer4):

you're adding and subtracting, so there's no way you can invent a new exponent, such as \(x^3\)...

OpenStudy (anonymous):

ohhh sorry...I'm really bad at math...I'm a writer...this would explain a lot I'm sure

OpenStudy (anonymous):

(surprising things? good things I hope? and not my math fail?)

OpenStudy (whpalmer4):

alas, I must dash your hopes on the rocks :-)

OpenStudy (anonymous):

x^4 + 5x^3 + 8x^3 + 40x^2 + 13x^2 +65x -x^2 -5x

OpenStudy (anonymous):

yes?

OpenStudy (whpalmer4):

distributive property: \[a(b+c) = a*b + a*c\]\[a(b-c) = a*b-a*c\] this means that \[-(b+c) = -1(b+c) = -1b -1c = -b-c\] the -( stuff ) construct trips up a lot of people, in my experience. Sometimes, writing it as the equivalent -1( stuff ) serves as enough of a visual reminder, but not always.

OpenStudy (whpalmer4):

Anyhow, we are just doing addition and subtraction of like terms here. We aren't able to create new exponents, so you only get the same exponents as you started with. \[x^2+8x+13 - (x^2+5x) = x^2+8x+13-1(x^2+5x)\]\[\qquad = x^2+8x+13 -1x^2 -1*5x\]\[\qquad = x^2 + 8x + 13 - x^2 -5x\]\[\qquad = x^2 - x^2 + 8x - 5x + 13\]\[\qquad = 3x+13\] any questions about that? We are NOT multiplying, just subtracting/adding

OpenStudy (whpalmer4):

If we draw this problem like a long division problem, this is what we have (so far): x ----------------- x + 5 | x^2 + 8x + 13 - (x^2 + 5x) | ---------------- 3x + 13

OpenStudy (anonymous):

nope, that makes perfect sense (I will attempt to not be such a fail...but no guarantees)

OpenStudy (whpalmer4):

does that make sense?

OpenStudy (anonymous):

it does

OpenStudy (whpalmer4):

Now, we repeat the process. We divide the first term of \(3x+13\) by the first term of \(x+5\); what is the result?

OpenStudy (whpalmer4):

\[\frac{3x}{x} = \]

OpenStudy (anonymous):

3

OpenStudy (whpalmer4):

Good. Now multiply 3 * the divisor; what do you get?

OpenStudy (anonymous):

3(x^2 + 8x + 13) = 3x^2 + 24 + 39

OpenStudy (anonymous):

24x

OpenStudy (anonymous):

my bad

OpenStudy (whpalmer4):

No, that's the dividend. The divisor is \(x+5\) \[3(x+5) = 3x+15\] x + 3 ----------------- x + 5 | x^2 + 8x + 13 - (x^2 + 5x) | ---------------- 3x + 13 - (3x + 15) ----------- -2 What does this mean?

OpenStudy (anonymous):

is this the remainder? and 3x +15 was the quotient?

OpenStudy (whpalmer4):

yes, no

OpenStudy (whpalmer4):

x+3 is quotient

OpenStudy (anonymous):

ohhhhhh, geez I feel dumb

OpenStudy (whpalmer4):

is that "geez I feel dumb" because you understand it now, or because you still don't understand it? :-)

OpenStudy (anonymous):

I think I'm gonna move on to english now...that at least is something I'm darn good at

OpenStudy (anonymous):

but thank you for the help! I became a fan because you know...your patience exceeds my imbecility

OpenStudy (anonymous):

or at least my sincere lack of competence in this field

OpenStudy (whpalmer4):

:-) it's funny, my sweetie says that in real life I have no patience whatsoever :-)

OpenStudy (anonymous):

you're biased with her, it's a whole different thing when you're dealing with a stranger

OpenStudy (whpalmer4):

there's another method of doing this called synthetic division. It only works for dividing by things without exponents, like \(x+5\), not like \(x^2+3x+5\) You write down all of the coefficients, including the "missing ones" (where you write down a 0 to represent a term that wasn't present, such as in \(x^2 + 16 = x^2 + 0x + 16\)) 1 8 13 -5 ------------ 1 you write down just the number from the divisor, and you write it with the opposite sign. so here I wrote -5 because our divisor is \(x+5\) finally, you drop the coefficient from the first column down below the line, as I did. Here's the drill. You take that right-most number from below the line, multiply it by the number "on the shelf" (-5, here), and write that underneath the number in the next column: 1 8 13 -5 -5 ------------ 1 now you add straight down: 1 8 13 -5 -5 ------------ 1 3 and repeat! 1 8 13 -5 -5 -15 ------------ 1 3 -2 we ran out of columns and have \(-2\) as our remainder. the "1 3" across the bottom is the coefficients of the quotient, so \(1x+3 = x+3\) just like we got up above. This goes very quickly, as you can see!

OpenStudy (whpalmer4):

if the the last number ends up being 0, there is no remainder.

OpenStudy (anonymous):

I remember doing this...I've taken Algebra 1, algebra 2, geometry, precalc and now I'm taking this math for college readiness

OpenStudy (whpalmer4):

For example: \[\frac{x^2+8x+15}{x+5}\] 1 8 15 -5 -5 -15 ----------- 1 3 0 so our answer is \[\frac{x^2+8x+15}{x+5} = x+3\]we can check by multiplying: \[(x+3)(x+5) = x(x+5) + 3(x+5) = x*x + 5*x + 3*x + 3*5 = x^2 + 5x + 3x + 15\]\[\qquad=x^2+8x+15\]just as it should be!

OpenStudy (anonymous):

it covers everrrrrryyyything...thank fully I only need one semester and not two since I have all my other math credits :D

OpenStudy (whpalmer4):

oops, ran off the edge of the "page" there, but no matter

OpenStudy (whpalmer4):

Well, I'll look forward to seeing you now and then ;-)

OpenStudy (anonymous):

oh yeah, you will...bugger this stupid credit...

OpenStudy (whpalmer4):

If you're good in English, you should try answering questions here in the English section. There are plenty of people who need help...

OpenStudy (whpalmer4):

Someone just tagged me with an English question, for example :-)

OpenStudy (anonymous):

I have actually :) it's real fun! and tagged you?

OpenStudy (whpalmer4):

I just tagged you on the question...